PFI-ZEKE SPECTROSCOPY OF ALUMINUM IMIDAZOLE AND PYRIMIDINE COMPLEXES

\underline{JUNG SUP LEE}, XU WANG, SERGIY KRASNOKUTSKI, and DONG-SHENG YANG, Department of Chemistry, University of Kentucky, Lexington, KY 40506-0055.

Aluminum complexes of imidazole (1,3-C_{3}H_{4}N_{2}) and pyrimidine (1,3-C_{4}H_{4}N_{2}) are produced by interactions between laser-ablated aluminum atoms with imidazole or pyrimidine vapor in supersonic molecular beams and studied by pulsed field ionization zero electron kinetic energy spectroscopy. The ionization energies of Al-imidazole and -pyrimidine are measured to be 35344(5) and 44182(6) cm^{−1}, and the Al^{+}-imidazole and -pyrimidine stretching frequencies are 308 and 262 cm^{−1}, respectively. In addition, an Al^{+}-imidazole in-plane bending mode is determined to be 125 cm^{−1}. Preliminary spectral analysis shows that aluminum atom or ion binds with single nitrogen of the ligand, and both complexes are in C_{s} symmetry.